# TiCoSS: Tightening the Coupling between Semantic Segmentation and Stereo Matching within A Joint Learning Framework

### v3

@@ III-C Hierarchical Deep Supervision
After tightening the coupling between semantic segmentation and stereo matching during the feature encoding stage, we turn our focus towards the feature decoding process. We first revisit the deep supervision strategies employed in SNE-RoadSeg+ [42] and UNet 3+ [55]. The former applies deep supervision to the decoded features with the highest resolution, whereas the latter achieves this on the deepest decoded features at each resolution. Despite the effectiveness of these two prior approaches in addressing challenges such as vanishing gradients and slow model convergence, the deep supervision strategies employed by them are not comprehensive (the former emphasizes enhancing local details, while the latter focuses on improving multi-scale feature consistency). Ideally, they should be used in conjunction to complement each other for improved results. Therefore, our proposed HDS strategy combines the strengths of both methods and demonstrates superior performance compared to each individually.
@@
A straightforward way to combine the strengths of these two methods is to apply deep supervision strategies simultaneously to both the main and side branches, enabling the network to leverage features from shallow layers (containing rich local details) and deep layers (being semantically strong). However, this method can not fully exploit the potential complementarity between the main and side auxiliary classifiers, since the decoded features are exclusively derived from adjacently-connected ones. Additionally, as discussed in Sect. III-B, decoded features in deep layers also lack informative spatial details which are typically contained in shallow layers, limiting the performance of DS strategies in our multi-task framework. Thus, to improve the interactions among auxiliary classifiers and provide local spatial details for decoded features in deep layers, we utilize the decoded, fused feature maps $\mathbf{F}_{1}^{F}$ (containing rich, fine-grained local spatial details that are essential for semantic segmentation) at the highest resolution to guide the feature decoding process in the side branches. Specifically, for the $l$-th auxiliary classifier within the side branch, we utilize a Feature Dynamic Alignment (FDA) block, which is composed of $l$ downsampling units to progressively align channel dimensions and spatial resolutions between a pair of features at different layers. Each downsampling unit comprises a $3\times 3$ convolutional layer with a stride of 2, followed by a batch normalization (BN) layer and a rectified linear unit (ReLU) activation layer. Compared to the simple max-pooling operation which may disrupt the original feature representation, our FDA achieves a smoother feature alignment. The feature maps obtained by downsampling $\mathbf{F}_{1}^{F}$ are then concatenated with the deepest decoded features at the corresponding layers. This downsampling output not only guides the decoding process but also serves as the input for the subsequent downsampling unit, thereby preserving fine-grained local details to the greatest extent possible. Since the outputs of the side branches are obtained by directly upsampling the deepest features at each layer, this strategy does not significantly impact training efficiency and memory usage. Moreover, the auxiliary classifiers of the main and side branches collaboratively provide additional pathways for gradients to propagate more efficiently from the output layers to their corresponding layers, thereby accelerating the convergence of our model.
@@ III-D Coupling Tightening Loss for Multi-task Joint Learning
$\mathcal{L}_{CT}=\mathcal{L}_{DIA}+\mathcal{L}_{DSCC}+\mathcal{L}_{SCG}+\mathcal{L}_{SM},$ (12)
where the DIA loss $\mathcal{L}_{DIA}$ measures disparity inconsistency, the DSCC loss $\mathcal{L}_{DSCC}$ measures the segmentation consistency across outputs from all deep supervision branches, the SCG loss $\mathcal{L}_{SCG}$ denotes the Semantic Consistency-Guided (SCG) loss proposed in [16], and $\mathcal{L}_{SM}$, a prevalently used loss function to supervise the training of the stereo matching network, is detailed in [16].
@@ III-D1 Disparity Inconsistency-Aware Loss
In our previous study [16], the developed SCG loss focuses mainly on enforcing semantic consistency to reduce segmentation errors caused by occlusions, neglecting the fact that occlusions can also lead to inconsistent disparity estimations [56], which are unfortunately under-explored when designing the loss. Thus, we further incorporate the disparity inconsistency in it to form a more advanced loss function to train our TiCoSS. Specifically, we define a weight matrix $\mathbf{W}\in\mathbb{R}^{H\times W}$ drawing from the concept of Left-Right Consistency in stereo matching [11], where
@@
$\mathcal{L}_{DIA}=\alpha\sum_{i=1}^{n}\left\{-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}\left[\mathbf{W}^{N}(\mathbf{p})\,y_{k}(\boldsymbol{p})\,\mathrm{log}(\hat{y}_{k}(\boldsymbol{p}))\right]\right\},$ (15)
where $n$ denotes the number of deep supervision branches, $N$ represents the total number of pixels, $C$ denotes the total number of classes, $\hat{y}_{k}(\boldsymbol{p})$ represents the predicted probability of pixel $\boldsymbol{p}$ belonging to class $k$, $y_{k}(\boldsymbol{p})$ denotes the ground-truth label for $\boldsymbol{p}$ in class $k$, and the weight factor $\alpha$ is determined through extensive ablation studies, as detailed in Sect. IV-E.
@@ III-D2 Deep Supervision Consistency Constraint Loss
In prior research [42, 55], the relationship among prediction maps generated by different auxiliary classifiers was not considered, which may lead to semantic inconsistencies across scales. To address this issue, we draw inspiration from [57] to design the following DSCC loss, which utilizes the KL divergence to measure the prediction differences across scales:
$\mathcal{L}_{DSCC}=\beta\sum_{r=1}^{L}\sum_{\begin{subarray}{c}s=1\\ s\neq r\end{subarray}}^{L}\left[-\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{k}^{r}(\boldsymbol{p})\log\frac{\hat{y}_{k}^{r}(\boldsymbol{p})}{\hat{y}_{k}^{s}(\boldsymbol{p})}\right],$ (16)
where l denotes the classifier index and $\beta$ is a scaling factor determined through extensive ablation studies presented in Sect. IV-E. By minimizing the DSCC loss, we enhance the informativeness of features at different resolutions, thereby significantly improving semantic segmentation performance across scales.
@@
Figure 4: Qualitative experimental results obtained under adverse weather conditions on the vKITTI2 [32] dataset, achieved by the SoTA semantic segmentation networks.
Figure 5: Qualitative experimental results on the KITTI 2015 [33] dataset achieved by the SoTA stereo matching networks.
Figure 6: Qualitative experimental results on the vKITTI2 [32] dataset achieved by the SoTA stereo matching networks.
@@ IV Experiments
In this article, we conduct extensive experiments to evaluate the performance of our introduced TiCoSS both quantitatively and qualitatively. The following subsections detail the used datasets, experimental set-up, evaluation metrics, and the comprehensive evaluation of our proposed method.
@@ IV-A Datasets
Since our network training requires both semantic and disparity annotations, we utilize the following two public datasets to conduct the experiments:
  • The vKITTI2 [32] dataset contains virtual replicas (providing 15 semantic classes) of five sequences from the KITTI dataset. Dense ground-truth disparity maps are obtained through depth rendering using a virtual engine. Following [16], we employ 700 stereo image pairs, along with their semantic and disparity annotations, to evaluate the effectiveness and robustness of our proposed TiCoSS, where 500 pairs are utilized for model training and the remaining 200 pairs are used for model validation.
@@ IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 4090 GPU with a batch size of 1. We set the maximum disparity search range to 192 pixels. All images are cropped to 512 $\times$ 256 pixels before being processed by the network. We utilize the AdamW [58] optimizer for model training, with epsilon and weight decay set to $10^{-8}$ and $10^{-5}$, respectively. The initial learning rate is set to $2\times 10^{-4}$. Training lasts for 100,000 iterations on the vKITTI2 dataset and 20,000 iterations on the KITTI 2015 dataset. Standard data augmentation techniques are applied to improve model robustness.
@@ IV-C Evaluation Metrics
Following [16], we quantify the performance of semantic segmentation using seven metrics: (1) accuracy (Acc), (2) mean accuracy (mAcc), (3) precision (Pre), (4) recall (Rec), (5) mean F1-score (mFSc), (6) mean intersection over union (mIoU), and (7) frequency-weighted intersection over union (fwIoU) [59]. Moreover, we quantify the performance of stereo matching using two metrics: (1) average end-point error (EPE) and (2) percentage of error pixels (PEP) at tolerance levels of 1.0 and 3.0 pixels, respectively.
TABLE I: Quantitative comparisons of SoTA semantic segmentation networks on the KITTI 2015 [33] dataset. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. Though more recent joint learning frameworks exist, the source code (in PyTorch or TensorFlow) for these previous approaches is not publicly available.
Networks Publication Type Acc (%) $\uparrow$ mAcc (%) $\uparrow$ mIoU (%) $\uparrow$ fwIoU (%) $\uparrow$ Pre (%) $\uparrow$ Rec (%) $\uparrow$ mFSc (%) $\uparrow$
PSPNet [5] CVPR’17 Single-Modal 80.03 44.97 38.15 68.62 79.29 82.66 79.59
DeepLabv3+ [37] ECCV’18 82.33 50.15 42.79 72.43 83.85 87.18 84.59
BiSeNet V2 [60] IJCV’21 73.68 41.66 32.71 60.55 68.35 81.79 72.37
Segmenter [38] ICCV’21 84.53 50.77 43.63 74.72 82.99 87.15 84.41
SegFormer [7] NeurIPS’21 88.28 59.23 51.39 80.53 87.15 90.85 88.46
Mask2Former [39] CVPR’22 84.35 54.33 45.87 75.56 84.74 89.12 85.92
MFNet [40] IROS’17 Feature-Fusion 81.02 48.13 40.70 70.42 82.85 85.73 82.36
RTFNet [41] RAL’19 71.61 39.26 30.35 57.98 69.52 85.16 74.28
SNE-RoadSeg [15] ECCV’20 79.46 51.91 41.56 69.22 81.45 87.05 82.91
OFF-Net [43] ICRA’22 75.84 40.13 33.13 64.02 77.48 72.19 70.62
RoadFormer [14] TIV’24 90.05 62.34 55.13 83.40 91.65 91.39 91.11
DispSegNet [20] RA-L’18 Joint Learning 88.03 62.54 53.44 80.90 88.22 92.53 89.68
DSNet[10] ICRA’19 89.48 64.44 52.83 81.08 89.41 93.25 90.73
S3M-Net [16] TIV’24 90.66 65.90 57.80 84.53 90.85 93.55 91.80
TiCoSS (Ours) / 91.90 71.97 63.63 86.68 92.43 94.10 92.90
@@
Networks Publication Type Acc (%) $\uparrow$ mAcc (%) $\uparrow$ mIoU (%) $\uparrow$ fwIoU (%) $\uparrow$ Pre (%) $\uparrow$ Rec (%) $\uparrow$ mFSc (%) $\uparrow$
PSPNet [5] CVPR’17 Single-Modal 76.26 53.53 44.81 69.30 81.55 79.68 75.38
DeepLabv3+ [37] ECCV’18 92.19 63.15 56.90 87.15 89.00 92.71 90.16
BiSeNet V2 [60] IJCV’21 81.77 51.07 44.45 74.71 83.23 82.19 80.67
Segmenter [38] ICCV’21 90.39 60.33 52.99 83.47 88.05 87.89 87.70
SegFormer [7] NeurIPS’21 94.75 70.56 64.98 90.49 93.57 93.62 93.46
Mask2Former [39] CVPR’22 89.29 64.58 57.14 83.84 90.75 87.23 87.19
MFNet [40] IROS’17 Feature-Fusion 76.22 51.50 43.41 68.82 82.46 78.65 73.80
RTFNet [41] RAL’19 85.22 49.47 42.59 77.69 83.74 89.17 84.41
SNE-RoadSeg [15] ECCV’20 83.64 60.85 52.56 75.14 83.44 81.66 77.77
OFF-Net [43] ICRA’22 90.84 61.51 55.27 84.69 89.24 86.71 86.15
RoadFormer [14] TIV’24 97.54 86.58 80.83 95.34 96.99 96.86 96.91
DispSegNet [20] RA-L’18 Joint Learning 97.15 82.63 80.24 95.21 97.39 97.66 97.46
DSNet[10] ICRA’19 95.71 78.71 77.47 93.24 97.20 96.66 96.81
S3M-Net [16] TIV’24 98.32 88.24 84.18 96.98 98.37 98.28 98.31
TiCoSS (Ours) / 98.69 91.66 88.46 97.58 98.55 98.67 98.57
TABLE III: Quantitative comparisons of semantic segmentation networks on the KITTI 2015 [33] dataset, evaluated using the mmsegmentation framework [61] . The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance.
Methods mIoU (%) $\uparrow$ mAcc (%) $\uparrow$
S3M-Net [16] 36.01 41.81
RoadFormer [14] 41.55 49.09
TiCoSS (Ours) 44.91 52.99
TABLE IV: Comparisons of SoTA stereo matching network on the KITTI 2015 [33] and vKITTI2 [32] datasets. The symbol $\downarrow$ indicates that a lower value corresponds to better performance. The best results are shown in bold font.

### v4

@@ III-C Hierarchical Deep Supervision
After tightening the coupling between semantic segmentation and stereo matching during the feature encoding stage, we turn our focus towards the feature decoding process. We first revisit the deep supervision strategies employed in SNE-RoadSeg+ [25] and UNet 3+ [46]. The former applies deep supervision to the decoded features with the highest resolution, whereas the latter achieves this goal on the deepest decoded features at each resolution. Despite the effectiveness of these two prior approaches in addressing challenges such as vanishing gradients and slow model convergence, the deep supervision strategies employed by them are not comprehensive (the former emphasizes enhancing local details, while the latter focuses on improving consistency across multi-scale segmentation predictions). Ideally, they should be used in conjunction to complement each other for improved results. Therefore, our proposed HDS strategy combines the strengths of both methods and demonstrates superior performance compared to each individually.
@@
A straightforward way to combine the strengths of these two methods is to apply deep supervision strategies simultaneously to both the main and side branches, enabling the network to leverage features from shallow layers (containing rich local details) and deep layers (being semantically strong). However, this method can not fully exploit the potential complementarity between the main and side auxiliary classifiers, since the decoded features are exclusively derived from adjacently-connected ones. Additionally, as discussed in Sect. III-B, decoded features in deep layers also lack informative spatial details that are present in shallow layers, limiting the performance of DS strategies in our multi-task learning framework. Thus, to improve the interactions among auxiliary classifiers and provide local spatial details for decoded features in deep layers, we utilize the decoded, fused feature maps $\mathbf{F}_{1}^{F}$ (containing rich, fine-grained local spatial details that are essential for semantic segmentation) at the highest resolution to guide the feature decoding process in the side branches. Specifically, for the $l$-th auxiliary classifier within the side branch, we utilize a feature dynamic alignment (FDA) block, which is composed of $l$ downsampling units to progressively align channel dimensions and spatial resolutions between a pair of features at different layers. Each downsampling unit comprises a $3\times 3$ convolutional layer with a stride of 2, followed by a batch normalization (BN) layer and a rectified linear unit (ReLU) activation layer. Compared to the simple max-pooling operation which may disrupt the original feature representation, our FDA achieves a smoother feature alignment. The feature maps obtained by downsampling $\mathbf{F}_{1}^{F}$ are then concatenated with the deepest decoded features at the corresponding layers. This downsampling output not only guides the decoding process but also serves as the input for the subsequent downsampling unit, thereby preserving fine-grained local details to the greatest extent. Since the outputs of the side branches are obtained by directly upsampling the deepest features at each layer, this strategy does not significantly impact training efficiency and memory usage. Moreover, the auxiliary classifiers of the main and side branches collaboratively provide additional pathways for gradients to propagate more efficiently from the output layers to their corresponding layers, thereby accelerating the convergence of our model.
@@ III-D Coupling Tightening Loss for Multi-Task Joint Learning
$\mathcal{L}_{CT}={\alpha\mathcal{L}_{DIA}+\beta\mathcal{L}_{DSCC}}+{\mathcal{L}_{SCG}+\mathcal{L}_{SM}}.$ (12)
where the DIA loss $\mathcal{L}_{DIA}$ measures disparity inconsistency, the DSCC loss $\mathcal{L}_{DSCC}$ measures the segmentation consistency across outputs from all deep supervision branches, the SCG loss $\mathcal{L}_{SCG}$ denotes the semantic consistency-guided (SCG) loss proposed in the study [5], $\mathcal{L}_{SM}$, a prevalently used loss function to supervise the training of the stereo matching network, is detailed in the study [5], and the weight factors $\alpha$ and $\beta$ are determined through extensive ablation studies, as detailed in Sect. IV-E.
@@ III-D1 Disparity Inconsistency-Aware Loss
In the previous study [5], the developed SCG loss focuses mainly on enforcing semantic consistency to reduce segmentation errors caused by occlusions, neglecting the fact that occlusions can also lead to inconsistent disparity estimations, which are unfortunately under-explored when designing the loss. Thus, we further incorporate the disparity inconsistency in it to form a more advanced loss function to train our TiCoSS. Specifically, we define a weight matrix $\mathbf{W}\in\mathbb{R}^{H\times W}$, drawing from the concept of left-right consistency check in stereo matching, where
@@
$\mathcal{L}_{DIA}=\sum_{i=1}^{n}\left\{-\frac{1}{N}\sum_{j=1}^{N}\sum_{k=1}^{C}\left[\mathbf{W}^{N}(\mathbf{p})\,y_{k}(\boldsymbol{p})\,\mathrm{log}(\hat{y}_{k}(\boldsymbol{p}))\right]\right\},$ (15)
where $n$ denotes the number of deep supervision branches, $N$ represents the total number of pixels, $C$ denotes the total number of classes, $\hat{y}_{k}(\boldsymbol{p})$ represents the predicted probability of pixel $\boldsymbol{p}$ belonging to class $k$, and $y_{k}(\boldsymbol{p})$ denotes the ground-truth label for $\boldsymbol{p}$ in class $k$.
@@ III-D2 Deep Supervision Consistency Constraint Loss
In prior studies [25, 46], the relationship among prediction maps generated by different auxiliary classifiers was not considered, which may lead to semantic inconsistencies across scales. To address this issue, we draw inspiration from the study [47] to design the following DSCC loss, which utilizes the KL divergence to measure the prediction differences across scales:
$\mathcal{L}_{DSCC}=\sum_{r=1}^{L}\sum_{\begin{subarray}{c}s=1\\ s\neq r\end{subarray}}^{L}\left[-\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{k}^{r}(\boldsymbol{p})\log\frac{\hat{y}_{k}^{r}(\boldsymbol{p})}{\hat{y}_{k}^{s}(\boldsymbol{p})}\right],$ (16)
where L denotes the total number of auxiliary classifiers. Further theoretical analyses of the DSCC loss are provided in the supplement.
@@
Figure 4: Qualitative experimental results achieved by the SoTA semantic segmentation networks on the Cityscapes [29] dataset.
@@ IV Experiments
In this article, we conduct extensive experiments to evaluate the performance of our introduced TiCoSS both quantitatively and qualitatively. The following subsections detail the used datasets, experimental setup, evaluation metrics, and the comprehensive evaluation of our proposed method.
@@ IV-A Datasets
Since our network training requires both semantic and disparity annotations, we utilize the following three public datasets to conduct the experiments:
  • The vKITTI2 [27] dataset contains virtual replicas (providing 15 semantic classes) of five sequences from the KITTI dataset. Dense ground-truth disparity maps are obtained through depth rendering using a virtual engine. Following the study [5], we employ 700 stereo image pairs, along with their semantic and disparity annotations, to evaluate the effectiveness and robustness of our proposed TiCoSS, where 500 pairs are utilized for model training and the remaining 200 pairs are used for model validation.
@@
  • The Cityscapes [29] dataset is a widely used urban scene understanding dataset, containing 2,975 stereo images for model training and 500 stereo images for model validation, with well-annotated semantic annotations. It is noteworthy that the corresponding depth information is obtained using ViTAStereo [43], since depth ground truth is unavailable.
@@ IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 3090 GPU with a batch size of 1. We set the maximum disparity search range to 192 pixels. All images are cropped to 512 $\times$ 256 pixels before being processed by the network. We utilize the AdamW optimizer for model training, with epsilon and weight decay set to $10^{-8}$ and $10^{-5}$, respectively. The initial learning rate is set to $2\times 10^{-4}$. Training lasts for 100,000 iterations on the vKITTI2 dataset, 20,000 iterations on the KITTI 2015 dataset, and 50,000 iterations on the Cityscapes dataset. Standard data augmentation techniques are applied to improve model robustness.
@@ IV-C Evaluation Metrics
Following the study [5], we quantify the performance of semantic segmentation using seven metrics: (1) accuracy (Acc), (2) mean accuracy (mAcc), (3) precision (Pre), (4) recall (Rec), (5) mean F1-score (mFSc), (6) mean intersection over union (mIoU), and (7) frequency-weighted intersection over union (fwIoU). We calculate each of these metrics on a per-image basis before averaging them across the entire dataset. Moreover, we quantify the performance of stereo matching using two metrics: (1) average end-point error (EPE) and (2) percentage of error pixels (PEP) at tolerance levels of 1.0 and 3.0 pixels, respectively.
TABLE I: Quantitative comparisons of SoTA semantic segmentation networks on the KITTI 2015 [28] dataset. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance.
Networks Publication Type Acc (%) $\uparrow$ mAcc (%) $\uparrow$ mIoU (%) $\uparrow$ Pre (%) $\uparrow$ Rec (%) $\uparrow$ mFSc (%) $\uparrow$
DeepLabv3+ [30] ECCV’18 Single-Modal 82.33 50.15 42.79 83.85 87.18 84.59
BiSeNet V2 [48] IJCV’21 73.68 41.66 32.71 68.35 81.79 72.37
Segmenter [31] ICCV’21 84.53 50.77 43.63 82.99 87.15 84.41
SegFormer [7] NeurIPS’21 88.28 59.23 51.39 87.15 90.85 88.46
Mask2Former [32] CVPR’22 84.35 54.33 45.87 84.74 89.12 85.92
RTFNet [33] RAL’19 Feature-Fusion 71.61 39.26 30.35 69.52 85.16 74.28
SNE-RoadSeg [3] ECCV’20 79.46 51.91 41.56 81.45 87.05 82.91
OFF-Net [34] ICRA’22 75.84 40.13 33.13 77.48 72.19 70.62
RoadFormer [14] TIV’24 90.05 62.34 55.13 91.65 91.39 91.11
RoadFormer+ [23] TIV’24 91.35 66.29 57.69 91.24 92.07 91.47
DFormer [35] ICLR’24 90.59 69.01 58.18 91.20 93.58 91.96
DispSegNet [16] RA-L’18 Joint Learning 88.03 62.54 53.44 88.22 92.53 89.68
DSNet[9] ICRA’19 89.48 64.44 52.83 89.41 93.25 90.73
SG-RoadSeg [19] ICRA’24 86.51 61.10 52.02 88.46 80.10 84.96
S3M-Net [5] TIV’24 90.66 65.90 57.80 90.85 93.55 91.80
TiCoSS (Ours) TASE’25 91.90 71.97 63.63 92.43 94.10 92.90
@@
Networks Publication Type Acc (%) $\uparrow$ mAcc (%) $\uparrow$ mIoU (%) $\uparrow$ Pre (%) $\uparrow$ Rec (%) $\uparrow$ mFSc (%) $\uparrow$
DeepLabv3+ [30] ECCV’18 Single-Modal 92.19 63.15 56.90 89.00 92.71 90.16
BiSeNet V2 [48] IJCV’21 81.77 51.07 44.45 83.23 82.19 80.67
Segmenter [31] ICCV’21 90.39 60.33 52.99 88.05 87.89 87.70
SegFormer [7] NeurIPS’21 94.75 70.56 64.98 93.57 93.62 93.46
Mask2Former [32] CVPR’22 89.29 64.58 57.14 90.75 87.23 87.19
RTFNet [33] RAL’19 Feature-Fusion 85.22 49.47 42.59 83.74 89.17 84.41
SNE-RoadSeg [3] ECCV’20 83.64 60.85 52.56 83.44 81.66 77.77
OFF-Net [34] ICRA’22 90.84 61.51 55.27 89.24 86.71 86.15
RoadFormer [14] TIV’24 97.54 86.58 80.83 96.99 96.86 96.91
RoadFormer+ [23] TIV’24 98.45 89.79 84.03 96.36 96.94 97.77
DFormer [35] ICLR’24 97.79 89.06 85.54 96.41 94.67 95.58
DispSegNet [16] RA-L’18 Joint Learning 97.15 82.63 80.24 97.39 97.66 97.46
DSNet[9] ICRA’19 95.71 78.71 77.47 97.20 96.66 96.81
SG-RoadSeg [19] ICRA’24 95.17 85.57 80.91 86.10 88.50 86.31
S3M-Net [5] TIV’24 98.32 88.24 84.18 98.37 98.28 98.31
TiCoSS (Ours) TASE’25 98.69 91.66 88.46 98.55 98.67 98.57
TABLE III: Quantitative comparisons of SoTA semantic segmentation networks on the Cityscapes [29] dataset. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance. Values marked with “-” denote that the corresponding metrics were not reported in the original paper.
Networks Publication Type Acc (%) $\uparrow$ mAcc (%) $\uparrow$ mIoU (%) $\uparrow$ Pre (%) $\uparrow$ Rec (%) $\uparrow$ mFSc (%) $\uparrow$
DeepLabv3+ [30] ECCV’18 Single-Modal 84.11 62.75 50.15 71.12 57.34 59.96
BiSeNet V2 [48] IJCV’21 85.61 61.16 49.65 71.00 57.61 62.50
Segmenter [31] ICCV’21 90.19 79.87 62.76 76.90 80.99 77.19
SegFormer [7] NeurIPS’21 88.12 73.79 62.48 75.50 78.81 71.02
Mask2Former [32] CVPR’22 87.97 78.80 64.29 75.01 78.26 74.71
RTFNet [33] RAL’19 Feature-Fusion 88.41 67.16 50.19 69.41 70.07 67.40
SNE-RoadSeg [3] ECCV’20 86.60 75.47 63.01 70.56 74.19 75.58
OFF-Net [34] ICRA’22 76.88 60.03 43.11 70.68 71.01 69.65
RoadFormer [14] TIV’24 87.11 75.49 62.20 74.49 76.84 75.50
RoadFormer+ [23] TIV’24 90.40 78.86 64.18 77.19 80.06 76.60
DFormer [35] ICLR’24 91.37 80.16 65.59 76.60 79.65 77.41
DispSegNet [16] RA-L’18 Joint Learning 84.07 62.18 56.10 70.00 60.92 61.16
DSNet [9] ICRA’19 84.01 62.25 52.79 69.56 63.74 61.26
MENet [12] TITS’21 93.39 69.53 61.50 - - -
SG-RoadSeg [19] ICRA’24 85.80 66.95 54.18 74.10 67.92 70.10
S3M-Net [5] TIV’24 88.47 77.30 62.59 75.20 77.48 76.30
TiCoSS (Ours) TASE’25 90.70 81.76 68.36 78.43 81.76 79.74
TABLE IV: Quantitative comparisons between TiCoSS and the baseline S3M-Net, evaluated using the mmsegmentation framework. The best results are shown in bold font. The symbol $\uparrow$ indicates that a higher value corresponds to better performance.
Methods mIoU (%) $\uparrow$
vKITTI2 KITTI 2015 Cityscapes
S3M-Net [5] 84.00 41.54 55.40
TiCoSS (Ours) 87.94 47.66 62.16
TABLE V: Comparisons of SoTA stereo matching network on the KITTI 2015 [28] and vKITTI2 [27] datasets. The symbol $\downarrow$ indicates that a lower value corresponds to better performance. The best results are shown in bold font.